OpenStudy (anonymous):

A hollow cylinder, open at one end, is constructed of thin sheet metal. The total external surface area of the cylinder is 192pi cm^2. The cylinder has a radius of r cm and a height of h cm. (i) express h in terms of r and show that the volume, Vcm^3 of the cylinder is given by V=pi/2(192r-r^3).

OpenStudy (anonymous):

i give up... sam's latex speed is even faster than my typing!

sam (.sam.):

Thats area \[A=\pi r^{2}+2\pi r h=192\pi\]

OpenStudy (anonymous):

Yes, I thought so.. I was just about to ask :)

sam (.sam.):

Find h from \[A=πr^{2}+2πrh=192π\]

OpenStudy (anonymous):

Oh. Ok.

OpenStudy (anonymous):

192pi-pi r^2 over 2 pi r =h

sam (.sam.):

h should be \[h=\frac{192-r^{2}}{2r}\]

OpenStudy (anonymous):

Yes

sam (.sam.):

\[V = πr^{2}h\] \[V = πr^{2}\frac{192-r^{2}}{2r}\]

OpenStudy (anonymous):

Ok. Thanks. There's a second part of the question. Should I post it here?

sam (.sam.):

post here

OpenStudy (anonymous):

Given that r can vary, (ii) find the value of r for which V has a stationary value.

OpenStudy (anonymous):

I mean V=0?

OpenStudy (anonymous):

Is that what I have to do?

sam (.sam.):

you need to differentiate V then set dV/dr=0

OpenStudy (anonymous):

So, how do I do that?

sam (.sam.):

\[V=πr^{2}h\] \[\frac{dV}{dr}=~~~96\pi-\frac{3\pi}{2}r^{2}~~=0~\] \[96\pi-\frac{3\pi}{2}r^{2}~~=0\] \[96\pi=\frac{3\pi}{2}r^{2}~~\] \[\frac{3}{2}r^{2}=96\] \[r^{2}=64\] r=8

OpenStudy (anonymous):

Wait, when you differentiate 96pi, doesn't that equal 0?

sam (.sam.):

\[V=\frac{\pi}{2}(192r-r^3)\] \[V=96\pi r-\frac{\pi r^{3}}{2}\] \[\frac{dV}{dr}=96\pi -\frac{3\pi r^{2}}{2}\]

OpenStudy (anonymous):

Oh, Ok. Silly me!

OpenStudy (anonymous):

The last part of the question says find this stationary value and determine whether it is maximum or minimum? Do I differentiate it again?

sam (.sam.):

r=8, substitute into \[V=96\pi r- \frac{\pi r^{3}}{2}\] \[V=512\pi\]

sam (.sam.):

for max or min, differentiate again

OpenStudy (anonymous):

would it be -3pi r = -3 pi times 8 = <0 = maximum point?

sam (.sam.):

yes

OpenStudy (anonymous):

Thank you :)!